# TI-DeepONet: Learnable Time Integration for Stable Long-Term Extrapolation

### v2

@@ 3.1 One-dimensional Burgers’ Equation
In contrast, the proposed TI-DeepONet variants, which embed numerical integration schemes into the learning framework, demonstrate superior extrapolation accuracy (see Figure 3(b)). TI-DeepONet with AB2/AM3 integration effectively stabilizes long-horizon predictions. TI(L)-DeepONet further improves on this by learning adaptive RK4 slope coefficients conditioned on the current state, enabling it to modulate time-stepping dynamically in response to the local solution behavior. This is particularly beneficial in stiff or nonlinear regimes, as reflected in the consistently lowest errors observed across all time steps. Figure A1 in SI presents iteration-wise learning of the RK4 slope coefficients, comparing it against the reference. An important point to note here is that the obtained optimized coefficients are compensating for the approximation error of the TI-DeepONet; hence, they may not necessarily adhere to the values obtained using the Taylor series expansion.
@@ 3.2 One-dimensional Korteweg-de Vries (KdV) Equation
where, $u$ is the amplitude of the wave, $\gamma$ and $\eta$ are some real-valued scalar parameters, $x$ and $t$ are the spatial and temporal dimensions, respectively. We generated 1,000 initial conditions to simulate time evolution, splitting them into 800 for training and 200 for testing.
@@
Figure 5 presents predictions from different frameworks on a representative sample. Both the full rollout and autoregressive DeepONet models fail to capture the dispersive, periodic dynamics of the solution. The full rollout approach, which treats time steps independently, lacks temporal continuity, while the autoregressive model suffers from compounding errors during recursive inference. Regarding the FNO variants, the full rollout FNO exhibits a noticeable phase shift due to its inability to capture the full spectrum of system frequencies; a consequence of being trained on a reduced temporal domain, thereby incurring significant errors throughout the entire temporal domain. The autoregressive FNO demonstrates particularly interesting behavior: while it achieves near-perfect predictions within the training domain, its performance deteriorates markedly upon entering the extrapolation regime. Here, autoregressive errors accumulate and manifest as spectral instability, characterized by frequency leakage. Specifically, the system’s frequencies become amplified at longer temporal horizons, resulting in erratic oscillations in the solution. Consequently, this approach also fails to reliably capture the inherent dispersive and periodic dynamics of the KdV PDE in the extrapolation domain. In contrast, TI-DeepONet and TI(L)-DeepONet closely match the ground truth, even in the extrapolation regime. The adaptive RK4 scheme in TI(L)-DeepONet yields similar accuracy as TI-DeepONet.
@@ 3.3 One-dimensional Kuramuto-Sivashinsky (KS) Equation
Here, $u$ represents the evolving scalar field with an initial profile $u_{0}$. The spatial domain $L$ is equipped with periodic boundary conditions. The nonlinear term $u\dfrac{\partial u}{\partial x}$ accounts for advection, while the dissipative term $\dfrac{\partial^{2}u}{\partial x^{2}}$ contributes to numerical stability, and the high-order dispersive term $\dfrac{\partial^{4}u}{\partial x^{4}}$ models the long-wavelength instability mechanisms that ultimately lead to chaotic dynamics. Further details regarding the data generation process are provided in SI C.3, where we adapt the data generation code from [li2021learning].
@@
Figure 6 presents the prediction errors incurred by the time-integrator–based neural operators compared to standard baselines, shown for both the training and extrapolation temporal domains for a representative sample. In the training regime, DON FR and FNO AR accurately recover the solution profile, whereas DON AR and FNO FR exhibit relatively higher errors. However, in the extrapolation domain, several interesting behaviors emerge for the standard NO baselines. DON FR, DON AR, and FNO FR incur substantial errors once predictions extend beyond the training horizon, with these errors growing progressively as extrapolation time increases. Notably, FNO AR produces highly noisy and fluctuating solution profiles, indicative of spectral instability or spectral bias artifacts. This behavior manifests as high-frequency leakage, where certain unresolved or weakly controlled high-frequency modes become excessively amplified, resulting in pronounced oscillations in the predicted solution. This phenomenon closely resembles the instability observed in FNO AR predictions for the 1D KdV system. Interestingly, while such behavior was observed only for the KdV (dispersive) and KS (chaotic) systems but not for Burgers (dissipative), we also found that the training of FNO AR was inherently unstable for certain random weight initializations. This suggests that systems exhibiting dynamics beyond standard diffusion may not be well-suited for the FNO autoregressive learning paradigm. In contrast, the TI-based variants demonstrate remarkably higher predictive accuracy in both the training and extrapolation temporal domains. Once again, TI(L)-DeepONet performs slightly better in capturing the unseen chaotic dynamics of the KS equation, which can be attributed to its solution-specific local adaptivity during the learning process.
@@
Figures 3(e) and (f) show the relative $L_{2}$ error growth over time for all frameworks. The autoregressive models, DON AR and FNO AR, are highly susceptible to error compounding under recursive rollout. Consequently, they begin accumulating errors within the initial timesteps, with these errors escalating rapidly in the extrapolation domain to reach values as high as 140% at the final timestep. FNO FR performs marginally better: while it still exhibits monotonically increasing error, the overall magnitude remains slightly lower, though still substantial in absolute terms. DON FR performs well in the training region due to DeepONet’s strong interpolation capabilities; however, its accuracy deteriorates rapidly upon entering the extrapolation domain, where errors grow steeply to reach levels comparable to FNO FR (approximately 80% at the final timestep). In stark contrast, the proposed TI-based variants leverage learned system dynamics by incorporating numerical time integrators within the learning framework. This coupling effectively suppresses error growth in the extrapolation region, even in the presence of chaotic dynamics. Consequently, these models not only achieve high accuracy in the training domain but also maintain relatively stable and controlled error growth throughout the extrapolation interval $t\in[15,30]$.
3.4 Two-dimensional Inviscid Burgers’ Equation
@@
where $u(x,y,t)$ is a scalar field and $\nu=10^{-4}$ is the kinematic viscosity. The initial condition, $u(x,y,0)=s(x,y)$, is sampled from 2D Gaussian random fields, with periodic boundary conditions applied in both spatial dimensions. Although the equation includes a finite viscosity term, its magnitude is sufficiently small ($\nu=10^{-4}$) as a result of which the system effectively behaves as inviscid, with shock formation expected at longer temporal horizons. This rich dynamical behavior, characterized by the development of sharp discontinuities and complex shock wave interactions, makes this problem an attractive benchmark for evaluating the performance of neural operator surrogates. SI C.4 contains comprehensive details of our data generation approach, which builds upon the code developed by [rosofsky2023applications].
@@
Figure 7: 2D Inviscid Burgers’ equation: Spatial error distribution across training ($t\in[0,0.33]$) and extrapolation ($t\in[0.33,1]$) regimes for all frameworks, illustrated with a representative sample.

### v3

@@ 3.1 One-dimensional Burgers’ Equation
In contrast, the proposed TI-DeepONet variants, which embed numerical integration schemes into the learning framework, demonstrate superior extrapolation accuracy (see Figures 3(a) and (b)). TI-DeepONet with AB2/AM3 integration (TI-DON/TI-DON AB) effectively stabilizes long-horizon predictions. TI(L)-DeepONet (TI(L)-DON) further improves on this by learning adaptive RK4 slope coefficients conditioned on the current state, enabling it to modulate time-stepping dynamically in response to the local solution behavior. This is particularly beneficial in stiff or nonlinear regimes, as reflected in the consistently lowest errors observed across all time steps. This observation is further reinforced in Figure 5, where the error distributions across test samples are comparable for both TI-DeepONet variants, with TI(L)-DeepONet exhibiting marginally smaller errors across a larger number of test samples. Among the baseline methods, the ranking in terms of test accuracy is: FNO AR >> DON FR >> FNO FR >> DON AR. Figure A1 in SI presents iteration-wise learning of the RK4 weighting coefficients, comparing them against the reference. An important point to note here is that the obtained optimized coefficients are compensating for the approximation error of the TI-DeepONet; hence, they may not necessarily adhere to the values obtained using the Taylor series expansion.
A detailed investigation into the behavior of the learned $\boldsymbol{\alpha}$ coefficients is presented in Figure 6. One can observe that the optimally learned $\alpha_{i}$’s differ significantly from the classical RK4 weighting coefficients $\{\frac{1}{6},\frac{2}{6},\frac{2}{6},\frac{1}{6}\}$. Notably, a higher emphasis is placed on the weight corresponding to the last computed slope $k_{4}$, with $\alpha_{4}$ exhibiting values closer to 1.0 compared to $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$. We hypothesize that this behavior arises because the slopes $k_{i}$ are computed using a neural network approximation of the time derivative, which inherently carries some approximation error. The classical RK4 weighting coefficients are derived under the assumption of exact slope evaluations to achieve 4th-order truncation error. However, when the slopes themselves are approximate, the optimal weighting may shift from minimizing truncation error to compensating for approximation error in the predicted time derivative. The auxiliary network learns coefficients $\boldsymbol{\alpha}$ that likely reflect a trade-off minimizing the overall prediction error for the specific dynamics being modeled. For the 1D Burgers’ equation, the network appears to discover that assigning higher weight to $k_{4}$, which is evaluated at the most advanced intermediate state $u^{n}+\Delta t\cdot k_{3}$, closest to the target $u^{n+1}$ yields more accurate predictions. This suggests that for smooth dynamics, the slope evaluation nearest to the target state may provide the most reliable estimate of the integrated dynamics over the timestep. A rigorous theoretical analysis of the order of accuracy and convergence properties of the learned adaptive weighting coefficients remains an important direction for future work.
@@ 3.2 One-dimensional Korteweg-de Vries (KdV) Equation
where, $u$ is the amplitude of the wave, $\gamma$ and $\eta$ are some real-valued scalar parameters, $x$ and $t$ are the spatial and temporal dimensions, respectively.
Data generation:
$u_{i}(x,0)=2k_{i}^{2}\,\text{sech}^{2}\left(k_{i}\left(\left(x+\frac{P}{2}-Pd_{i}\right)\%P-\frac{P}{2}\right)\right)$ (7)
where sech is the hyperbolic secant, $P$ is the period in space, % is the modulo operator, and $i=\{1,2\}$. The coefficients $k_{i}$ and $d_{i}$, which determine the height and location of the soliton peak, respectively, are independently and uniformly sampled from the intervals $k_{i}\in[0.3,0.7]$ and $d_{i}\in[0,1]$ for each soliton. The spatial period is set to $P=10$. The dataset is curated by generating $N=1000$ distinct initial conditions and their temporal evolution is modeled using the midpoint method. The 1D spatial domain, $\Omega=[0,10]$, is discretized into 100 uniformly spaced grid points, resulting in a spatial resolution of $\Delta x=0.1$. For the temporal domain, we consider a uniform grid of 201 points spanning the time domain $t\in[0,5]$, which yields $\Delta t=0.025$. For further details regarding the problem definition, the reader is referred to Michałowska et al. (2024).
Data preparation:
Training architecture:
Tables A4 and A5 present the network architecture details, including the number of layers, neurons per layer, activation functions, total epochs to convergence, batch size, and Fourier modes (where applicable). For the TI(L)-DeepONet framework, we employ a secondary feedforward network to predict the four RK4 slope coefficients. This network consists of two hidden layers with 64 and 48 neurons, respectively, followed by an output layer with four neurons. The hidden layers use the tanh activation function, while a softmax activation is applied at the output layer to ensure that the predicted coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 epochs. A similar learning rate schedule was used for training the auxiliary network tasked with learning the RK4 slope coefficients. The same learning rate schedule with identical hyperparameters was employed for FNO FR training. For FNO AR, an exponential decay learning rate schedule was also used, but with an initial learning rate of $10^{-3}$ that decayed by a factor of 0.96 every 2000 steps.
@@
Figure 8: 1D KdV Equation: Distribution of relative $L_{2}$ errors across all test samples for all frameworks, averaged over the entire spatiotemporal domain.
Figure 7 presents predictions from different frameworks on a representative sample. Both DON FR and DON AR models fail to capture the dispersive, periodic dynamics of the solution. The full rollout approach, which treats time steps independently, lacks temporal continuity, while the autoregressive model suffers from compounding errors during recursive inference. Regarding the FNO variants, FNO FR exhibits a noticeable phase shift due to its inability to capture the full spectrum of system frequencies; a consequence of being trained on a reduced temporal domain, thereby incurring significant errors throughout the entire temporal domain. FNO AR demonstrates particularly interesting behavior: while it achieves near-perfect predictions within the training domain, its performance deteriorates markedly upon entering the extrapolation regime. Here, autoregressive errors accumulate and manifest as spectral instability, characterized by frequency leakage. Specifically, the system’s frequencies become amplified at longer temporal horizons, resulting in erratic oscillations in the solution. Consequently, this approach also fails to reliably capture the inherent dispersive and periodic dynamics of the KdV PDE in the extrapolation domain. In contrast, TI-DON and TI(L)-DON closely match the ground truth, even in the extrapolation regime. The adaptive RK4 weighting scheme in TI(L)-DeepONet yields marginally improved accuracy compared to TI-DeepONet.
@@ 3.3 One-dimensional Kuramuto-Sivashinsky (KS) Equation
Here, $u$ represents the evolving scalar field with an initial profile $u_{0}$. The spatial domain $L$ is equipped with periodic boundary conditions. The nonlinear term $u\dfrac{\partial u}{\partial x}$ accounts for advection, while the dissipative term $\dfrac{\partial^{2}u}{\partial x^{2}}$ contributes to numerical stability, and the high-order dispersive term $\dfrac{\partial^{4}u}{\partial x^{4}}$ models the long-wavelength instability mechanisms that ultimately lead to chaotic dynamics.
Data generation:
We consider the 1D Kuramoto-Sivashinsky (KS) equation on a spatial domain $x\in[0,L]$ with $L=6\pi$, prescribed with periodic boundary conditions. The initial conditions $u(x,0)=s(x)$ are sampled from mean-zero Gaussian random fields (GRFs) with periodic boundary conditions. Specifically, we generate GRFs with covariance operator $\mathcal{C}=\sigma^{2}(-\Delta+\tau^{2}I)^{-\gamma}$, where $\gamma=1$, $\tau=2$, and $\sigma=1$. The spatial domain is discretized using $N_{x}=128$ grid points, and the temporal domain $t\in[0,30]$ is discretized with $N_{t}=300$ timesteps. The PDE is solved using the exponential time-differencing fourth-order Runge-Kutta (ETDRK4) method with a timestep of $\Delta t=0.1$. We generate a total of 3000 samples, each consisting of the initial condition and the corresponding spatiotemporal solution field $u(x,t)$. The dataset is partitioned using a train-test split ratio of 0.8, yielding $N_{\text{train}}=2400$ training samples and $N_{\text{test}}=600$ testing samples. The periodic boundary conditions ensure that the solution remains in the space of mean-zero periodic functions throughout the evolution, preserving the chaotic dynamics characteristic of the KS equation. For further details regarding the data generation procedure, the reader is referred to Li et al. (2021), from which the data generation code is adapted.
Data preparation:
Training architecture:
Tables A6 and A7 present the network architecture details, including the number of layers, neurons per layer, activation functions, total epochs to convergence, batch size, and Fourier modes (where applicable). All DeepONet variants employ Fourier feature encoding in the trunk network. For the full rollout DeepONet, both spatial and temporal coordinates $(x,t)$ are augmented with sinusoidal features computed as $\sin(\pi kx),\cos(\pi kx),\sin(\pi kt),\cos(\pi kt)]$ for $k\in\{1,2,\ldots,10\}$, expanding the input dimension from 2 to 42. For the autoregressive and TI-based variants, only the spatial coordinate $x$ is augmented with Fourier features $\sin(\pi kx),\cos(\pi kx)]$, expanding the input dimension from 1 to 21. For the TI(L)-DeepONet framework, we employ an auxiliary feedforward network to predict the four RK4 slope coefficients. This network consists of two hidden layers with 64 neurons each, followed by an output layer with four neurons. The hidden layers utilize the tanh activation function, while a softmax activation is applied at the output layer to ensure the predicted coefficients sum to one. All DeepONet variants (DON FR, DON AR, TI-DON, and TI(L)-DON) are trained using the AdamW optimizer with an initial learning rate of $10^{-3}$, exponentially decayed by a factor of 0.95 every 5000 epochs, and a weight decay of $10^{-4}$. For TI(L)-DeepONet specifically, the auxiliary RK4 network is trained separately using the Adam optimizer with an initial learning rate of $2\times 10^{-3}$ and the same exponential decay schedule. The FNO variants employ the Adam optimizer with an initial learning rate of $10^{-3}$, exponentially decayed by a factor of 0.96 every 2000 steps for both FNO FR and FNO AR configurations.
@@
Figure 10: 1D KS Equation: Distribution of relative $L_{2}$ errors across all test samples for all frameworks, averaged over the entire spatiotemporal domain.
@@
Figure 9 presents the prediction errors incurred by the TI-based neural operators compared to standard baselines, shown for both the training and extrapolation temporal domains for a representative sample. In the training regime, DON FR and FNO AR accurately recover the solution profile, whereas DON AR and FNO FR exhibit relatively higher errors. However, in the extrapolation domain, several interesting behaviors emerge for the standard NO baselines. DON FR, DON AR, and FNO FR incur substantial errors once predictions extend beyond the training horizon, with these errors growing progressively as extrapolation time increases. Notably, FNO AR produces highly noisy and fluctuating solution profiles, indicative of spectral instability or spectral bias artifacts. This behavior manifests as high-frequency leakage, where certain unresolved or weakly controlled high-frequency modes become excessively amplified, resulting in pronounced oscillations in the predicted solution. This phenomenon closely resembles the instability observed in FNO AR predictions for the 1D KdV system. Interestingly, while such behavior was observed only for the KdV (dispersive) and KS (chaotic) systems but not for Burgers (dissipative), we also found that the training of FNO AR was inherently unstable for certain random weight initializations. This suggests that systems exhibiting dynamics beyond standard diffusion may not be well-suited for the FNO autoregressive learning paradigm. In contrast, the TI-based variants demonstrate remarkably higher predictive accuracy in both the training and extrapolation temporal domains. Once again, TI(L)-DON performs slightly better in capturing the unseen chaotic dynamics of the KS equation, which can be attributed to its solution-specific local adaptivity during the learning process.
Figure 11: 1D KS Equation: Distribution of the optimally learned $\alpha_{i}~\forall i\in\{1,2,3,4\}$ for the RK4 slope coefficients.
@@
Figures 3(e) and (f) show the relative $L_{2}$ error growth over time for all frameworks. The autoregressive models, DON AR and FNO AR, are highly susceptible to error compounding under recursive rollout. Consequently, they begin accumulating errors within the initial timesteps, with these errors escalating rapidly in the extrapolation domain to reach values as high as 140% at the final timestep. FNO FR performs marginally better: while it still exhibits monotonically increasing error, the overall magnitude remains slightly lower, though still substantial in absolute terms. DON FR performs well in the training region due to DeepONet’s strong interpolation capabilities; however, its accuracy deteriorates rapidly upon entering the extrapolation domain, where errors grow steeply to reach levels comparable to FNO FR (approximately 80% at the final timestep). In stark contrast, the proposed TI-based variants leverage learned system dynamics by incorporating numerical time integrators within the learning framework. This coupling effectively suppresses error growth in the extrapolation region, even in the presence of chaotic dynamics. Consequently, these models not only achieve high accuracy in the training interval but also maintain relatively stable and controlled error growth throughout the extrapolation interval $t\in[15,30]$. Between the TI-based variants, TI(L)-DON marginally outperforms TI-DON, as evidenced in Table 1 and Figure 3(e). The error distributions across test samples, shown in Figure 10, remain largely similar for both variants, with TI(L)-DON exhibiting slightly better performance characterized by a higher density of samples with errors under 2% compared to TI-DON. This indicates that the solution-specific local adaptivity of TI(L)-DON is effective in capturing the underlying stiff, chaotic dynamics of the KS equation. On the contrary, the test performance of other NO baselines consistent with the previous observations remains as: DON FR > FNO FR > FNO AR > DON AR. In a similar spirit to the previous cases, we present the distribution of the learned weighting coefficients $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$, and $\alpha_{4}$ in Figure 11. Consistent with earlier observations, $\alpha_{4}$ predominantly takes large values close to 1, indicating that greater emphasis is placed on the slope $k_{4}$ for a majority of solution states. However, unlike the previous cases, a subset of solution states exhibits lower values of $\alpha_{4}$ accompanied by higher values of $\alpha_{1}$, suggesting that for certain solution configurations, the network learns to place greater weight on the initial slope $k_{1}$. This variability in the learned coefficients reflects the complex, chaotic nature of the KS dynamics, where different solution states may benefit from different weighting strategies within the adaptive RK4 scheme.
3.4 Two-dimensional Burgers’ Equation
@@
where $u(x,y,t)$ is a scalar field and $\nu=10^{-4}$ is the kinematic viscosity. The initial condition, $u(x,y,0)=s(x,y)$, is sampled from 2D Gaussian random fields, with periodic boundary conditions applied in both spatial dimensions. Although the equation includes a finite viscosity term, its magnitude is sufficiently small ($\nu=10^{-4}$) as a result of which the system approximately behaves as inviscid, with shock formation possible at longer temporal horizons. This rich dynamical behavior, characterized by the development of sharp discontinuities and complex shock wave interactions, makes this problem an attractive benchmark for evaluating the performance of neural operator surrogates. The following subsections contain comprehensive details of our data generation approach, which builds upon the code developed by Rosofsky et al. (2023).
Data generation:
For the 2D Burgers’ equation, we consider periodic boundary conditions on a unit square domain $(x,y)\in[0,1]^{2}$. The initial conditions $u(x,y,0)=s(x,y)$ are sampled from periodic Matérn-type Gaussian random fields with length scale $l=0.1$ and standard deviation $\sigma=0.2$. The spatial domain is discretized using a $64\times 64$ grid with $N_{x}=N_{y}=64$ points. The PDE is solved using a fourth-order Runge-Kutta scheme with timestep $\Delta t=10^{-4}$ up to final time $t=1.0$. Solutions are saved at intervals of $\Delta t_{\text{save}}=0.01$, resulting in 101 temporal snapshots. The kinematic viscosity is set to $\nu=10^{-4}$, which is sufficiently small to induce inviscid behavior while maintaining numerical stability. We generate a total of 1000 samples, each consisting of the initial condition and the corresponding spatiotemporal solution field $u(t,x,y)$ of size $101\times 64\times 64$.
Data preparation:
The dataset is partitioned using a train-test split ratio of 0.8, yielding $N_{\text{train}}=800$ training samples and $N_{\text{test}}=200$ testing samples. To evaluate extrapolation capabilities, we train only on the first third of the temporal domain, i.e., $t_{\text{train}}\in[0,0.33]$ (33 out of 101 time steps), and assess performance over the entire domain $t\in[0,1]$. This represents one of the most challenging extrapolation scenarios in our study, requiring $3\times$ temporal extrapolation with no exposure to shock formation during training. For full rollout methods, our aim is to learn a solution operator $\mathcal{G}_{\boldsymbol{\theta}}$ that maps initial conditions to the full solution: $u_{0}(x,y)\rightarrow u(t,x,y)$. The trunk network receives $(t_{\text{train}},x,y)$ coordinates as input, while the branch network processes the initial conditions. For FNO FR, the 3D Fourier transform is performed along all three dimensions $(t_{\text{train}},x,y)$, capturing the full spatiotemporal frequency content within the training domain. In contrast, autoregressive and TI-based methods use only spatial coordinates $(x,y)$ for the trunk network input. Their branch networks receive stacked solution states $u^{0}(x,y),u^{1}(x,y),\ldots,u^{33}(x,y)]$ from the training temporal window. FNO AR employs a 2D Fourier transform along the spatial dimensions $(x,y)$ only, consistent with its autoregressive nature to learn a one-step mapping between pairs of consecutive solution states $(u^{i}(x,y),u^{i+1}(x,y))$. Note that for the autoregressive/TI-based variants, explicit temporal information is absent and instead is assimilated in the batch dimension.
Training architecture
Tables LABEL:tab:2d_burgers_don-architecture and A9 presents the network architecture details, including the number of layers and neurons per layer, activation functions, total epochs to convergence, batch size and Fourier modes (if applicable). All DeepONet variants employ Fourier feature encoding in the trunk network. For DON FR, the spatial and temporal coordinates $(x,y,t)$ are augmented with sinusoidal features computed as $\sin(\pi kx),\cos(\pi kx),\sin(\pi ky),\cos(\pi ky),\sin(\pi kt),\cos(\pi kt)]$ for $k\in\{1,2,\ldots,10\}$, expanding the input dimension from 3 to 63. For the autoregressive and TI-based variants, only the spatial coordinates $(x,y)$ are augmented with Fourier features $\sin(\pi kx),\cos(\pi kx),\sin(\pi ky),\cos(\pi ky)]$, expanding the input dimension from 2 to 42. In this case, for the TI(L)-DON architecture, we employ a similar combination of 2D convolutional blocks and MLP layers as used in the branch network to learn the RK4 slope coefficients. The convolutional blocks utilize the same combination of Conv2D and pooling layers as the branch, with the only difference being that the Conv2D layers have a reduced number of filters (32 instead of 64). This is followed by a flattening layer, after which the intermediate representation is passed through two fully connected hidden layers and finally to an output layer with 4 neurons that predicts the RK4 slope coefficients. The Conv2D layers use a ReLU activation function, the fully connected hidden layers use tanh, and the output layer applies a softmax activation to ensure that the predicted slope coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 steps. The auxiliary network responsible for predicting the learnable RK4 slope coefficients was trained with a similar exponential decay schedule, but with a higher initial learning rate of $2\times 10^{-3}$. For both FNO AR and FNO FR variants, we employ a similar exponential decay schedule as the DeepONet variants albeit with a decay factor of 0.96 every 2000 steps.
@@
Figure 12: 2D Burgers’ Equation: Spatial error distribution across training ($t\in[0,0.33]$) and extrapolation ($t\in[0.33,1]$) regimes for all frameworks, illustrated with a representative sample.
Figure 13: 2D Burgers’ Equation: Distribution of relative $L_{2}$ errors across all test samples for all frameworks, averaged over the entire spatiotemporal domain.
Figure 14: 2D Burgers’ Equation: Distribution of the optimally learned $\alpha_{i}~\forall i\in\{1,2,3,4\}$ for the RK4 slope coefficients.